OpenStudy (anonymous):

Find y' of x^2*(x-y)^2=x^2-y^2

OpenStudy (luigi0210):

Did you take the derivative of it yet?

OpenStudy (anonymous):

Yeah but I think I screwed up the left side somehow hahah

OpenStudy (luigi0210):

Hmm, put what you got and we'll see what we can do about it :P

OpenStudy (anonymous):

I have 2x(x-y)^2*x^2*2(x-y)(1-y) = 2x-2y*y

OpenStudy (anonymous):

Wait need to put y'

OpenStudy (anonymous):

2x(x-y)^2*x^2*2(x-y)(1-y')=2x-2y*y' ?

OpenStudy (solomonzelman):

yes, now you need to isolate the y'

OpenStudy (solomonzelman):

wait, there should be a plus in the middle on the left hand side

OpenStudy (loser66):

why don't we simplify it to get the simpler equation?

OpenStudy (luigi0210):

FOILing that looks like a lot of work xD

OpenStudy (solomonzelman):

x^2*(x-y)^2=x^2-y^2 `2x(x-y)^2` + `x^2 2(x-y)(1-y')` = 2x-2y*y'

OpenStudy (anonymous):

Yes you're right I put a * instead of a + I screwed that up

OpenStudy (solomonzelman):

expanding? maybe... I don't find a need in doing that. But that is just me

OpenStudy (loser66):

\(x^2(x-y)^2=x^2-y^2\\x^2(x-y)(x-y) =(x-y)(x+y)\\x^2(x-y) =x+y\)

OpenStudy (loser66):

then expand and isolate y, take derivative then. you will be ok

OpenStudy (solomonzelman):

yes, that is definitely an option:) agree...

OpenStudy (solomonzelman):

not that it is super hard to take the product rule at the beginning

OpenStudy (anonymous):

Awesome thanks for the help guys!

OpenStudy (solomonzelman):

yup.... I am sure you can isolate the y' (isolation should be something automatic, once you are in calc) even I can do it:)